OpenStudy (anonymous):

ALGEBRA 2 HELP (will give a medal)

OpenStudy (anonymous):

what is the simplest form of the expression sqrt3 - sqrt6 / sqrt3 + sqrt6

jimthompson5910 (jim_thompson5910):

They gave you \[\Large \frac{\sqrt{3}-\sqrt{6}}{\sqrt{3}+\sqrt{6}}\] right?

OpenStudy (anonymous):

yes. i have answer choices too, if that helps

jimthompson5910 (jim_thompson5910):

the given denominator is \(\Large \sqrt{3}+\sqrt{6}\) the conjugate is \(\Large \sqrt{3}-\sqrt{6}\) Notice how the + changed to - Multiply top and bottom by the conjugate of the denominator to get... \[\Large \frac{\sqrt{3}-\sqrt{6}}{\sqrt{3}+\sqrt{6}}\] \[\Large \frac{(\sqrt{3}-\sqrt{6}){\color{red}{*(\sqrt{3}-\sqrt{6})}}}{(\sqrt{3}+\sqrt{6}){\color{red}{*(\sqrt{3}-\sqrt{6})}}}\] In the denominator, you can use the rule (a-b)(a+b) = a^2 - b^2 to take a shortcut and you'll notice the radicals go away in the denominator I'll let you finish up

OpenStudy (anonymous):

So would I do (sqrt3-sqrt6)(sqrt3+sqrt6) = 3^2 - 6^2 ?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

you square "sqrt(3)" to get 3 you square "sqrt(6)" to get 6 ie \[\Large (\sqrt{3})^2-(\sqrt{6})^2 = 3 - 6 = -3\]

jimthompson5910 (jim_thompson5910):

the square and square root are opposite operations they undo and cancel each other out

jimthompson5910 (jim_thompson5910):

you now have \[\Large \frac{(\sqrt{3}-\sqrt{6})(\sqrt{3}-\sqrt{6})}{-3}\]

jimthompson5910 (jim_thompson5910):

see what happens when you expand out the numerator

OpenStudy (anonymous):

okay, would i multiply (sqrt3 - sqrt6)(sqrt3 - sqrt6)?

jimthompson5910 (jim_thompson5910):

yes, multiply that out

OpenStudy (anonymous):

do i want to keep it as a square root?

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

never mind. i put it in my calculator and wound up with 0.514718625761. but i dont think that's right

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):